# dInfer: An Efficient Inference Framework for Diffusion Language Models

### v1

@@ 1 Introduction
Over the past year, diffusion-based large language models (dLLMs) have gained increasing attention in both academia and industry. Unlike conventional autoregressive (AR) models that generate tokens sequentially, dLLMs refine entire sequences in parallel through iterative denoising. This intrinsic parallelism opens new opportunities for faster decoding and enables better utilization of GPU hardware. Combined with rapid algorithmic progress, these properties make dLLMs a compelling alternative to AR LLMs. Recent work has shown that models such as LLaDA(-MoE) Nie et al. (2025), Zhu et al. (2025) can reach performance levels comparable to strong AR baselines, including Llama Grattafiori et al. (2024) and Qwen Yang et al. (2024).
@@
To address these challenges, we propose dInfer, an efficient and extensible inference framework for dLLMs. dInfer modularizes inference into four components—model, diffusion iteration manager, decoding strategy, and KV-cache management—and provides well-designed APIs for flexible combinations of algorithms in each component. It supports multiple dLLM variants, including LLaDA Nie et al. (2025), LLaDA-MoE Zhu et al. (2025), and LLaDA-MoE-TD (Section C.1). dInfer introduces an iteration smoothing algorithm for smoother denoising, hierarchical and credit decoding for enhanced parallel decoding, and a vicinity refresh strategy for KV-cache management to mitigate cache staleness.
@@
We evaluate inference efficiency using tokens per second (TPS) per sequence, shown in Figure 1. On HumanEval, dInfer achieves over 1,100 TPS at batch size 1, and averages more than 800 TPS across six benchmarks on a single node with $8\times$ H800 GPUs. Compared to Fast-dLLM Wu et al. (2025), dInfer delivers more than a $10\times$ speedup while maintaining accuracy; on LLaDA-MoE it provides a $2-3\times$ speedup over QWen2.5-3B on vLLM with comparable quality.
@@ 2 Framework Design
A key advantage of dLLMs lies in their ability to perform parallel decoding within each iteration. To fully exploit this capability, advances are required in model design, diffusion iteration strategies, and decoding algorithms. At the same time, dLLMs introduce unique computational challenges. Unlike AR models—where previously computed keys and values can be cached—dLLMs employ bidirectional attention, meaning that decoding a single token can affect the representations of all tokens in the sequence. This makes straightforward KV-cache reuse infeasible and necessitates specialized cache management for efficient inference.
@@ 2.1 Diffusion Iteration Manager
dInfer currently implements two algorithms in this component. The block-wise diffusion iteration algorithm performs decoding in fixed-size spans and serves as a baseline (Algorithm 1). The iteration smoothing algorithm improves upon this by retaining token representations from the previous iteration and fusing them with the next iteration’s embeddings. This enables cross-iteration information flow, enriches contextual cues, and empirically boosts token confidence while mitigating the performance degradation typically caused by KV-cache deployment. A detailed description is provided in Appendix A.1.
@@ 2.2 Decoding Strategy
  • Threshold decoding (from Fast-dLLM Wu et al. (2025)): commits tokens whose confidence exceeds a preset threshold.
@@ 2.3 KV-cache management
Earlier approaches introduced training-free strategies such as block-wise caching and Dual Cache Wu et al. (2025), which reuse KV states for decoded tokens or suffixes of masked tokens. However, these methods treat cached states as static, neglecting updates from newly decoded tokens, which often degrades accuracy.
@@ 2.5 An example of the orchestration of the algorithms in dInfer
Figure 3 shows how the algorithms interact within dInfer. In each iteration, the framework tries to identify the [MASK] tokens in the active decoding block. The sequence is first embedded (potentially leveraging context from previous iterations), followed by a forward pass through the model using tensor/expert parallelism to produce logits. During the forward pass, the previous KV cache that does not hit the vicinity refresh strategy will remain unchanged and be reused. After getting the logits, the hierarchical/credit decoding algorithm will decide which [MASK] tokens to decode and predict their identities. In addition, the iteration smoothing algorithm will retain the logit-weighted embeddings of the sequence and incorporate them as part of the embedding for the next iteration, ensuring continuity across steps. Together, these components enable dInfer to achieve both efficiency and stability in dLLM inference.
@@ 3 Implementation Details
dInfer builds on vLLM’s backend to exploit two complementary forms of parallelism. Tensor parallelism is applied to the linear layers preceding attention modules, distributing dense computations efficiently across multiple GPUs. Expert parallelism is applied to the LLaDA-MoE model and is effective even at a batch size of one—unlike in AR models, where expert parallelism typically requires large batch sizes. By combining tensor and expert parallelism, dInfer achieves more than a 100% improvement in inference efficiency.
@@
Diffusion iterations can suffer from CUDA stream bubbles—idle gaps of consecutive kernel launches between diffusion iterations—which waste GPU cycles and reduce throughput. To address this, dInfer applies a loop unrolling strategy that allows Python to launch CUDA kernels continuously without being blocked by stream synchronization. This reduces launch latency, keeps GPU pipelines fully occupied, and boosts iteration efficiency by about 5-10%.
We also introduce an early termination mechanism for block-wise decoding. Once an end-of-sequence (EOS) token is generated within a block, subsequent decoding steps on the remaining blocks become redundant. dInfer therefore halts the diffusion loop and fills all remaining blocks with EOS, avoiding unnecessary computation. This optimization improves inference efficiency by 15-40%.
@@ 4.1 Datasets and Configurations
All experiments are conducted on a server equipped with $8\times$ NVIDIA H800 GPUs, with PyTorch 2.9.0.dev20250831 and vLLM 0.10.1. We use a batch size of one, a generation length of 1024 and a block size of 64 for all experiments.
@@ 4.2 Performance
As shown in Table 1, it can be clearly observed that LLaDA-MoE have comparable or higher performance than Qwen2.5-3B over six diverse datasets. Under the ”Without KV Cache” setting, dInfer achieves an average accuracy of 54.33, which is higher than Fast-dLLM and is comparable to the performance of QWen2.5-3B in vLLM and LLaDA-MoE reported in its paper Zhu et al. (2025), while achieving $6.5\times$ speedup over Fast-dLLM. When KV cache is enabled, dInfer achieves higher accuracy than Fast-dLLM (53.96 vs. 52.15) and delivers $6\times$ speedup over Fast-dLLM. When achieving similar model performance, dInfer achieves over $10\times$ speedup (a TPS of 680.71) over Fast-dLLM (a TPS of 63.61) across the six benchmarks. dInfer is also $2.5\times$ faster than Qwen-2.5 3B (a TPS of 277.45) in vLLMs.
@@ 5 Conclusion
In this work, we presented dInfer, an efficient and extensible inference framework for dLLMs. By decomposing inference into modular components and incorporating optimizations such as hierarchical decoding, credit decoding, iteration smoothing, and vicinity KV-cache refresh, dInfer effectively addresses the key bottlenecks of high computation cost and limited parallel decoding efficiency of dLLMs. Extensive experiments on LLaDA-MoE demonstrate that dInfer achieves state-of-the-art throughput—exceeding 1,100 TPS on $8\timesH$800 GPUs—while maintaining output quality. We believe dInfer provides both a practical toolkit and a standardized platform to accelerate research and development in the rapidly growing field of dLLMs.
@@ A.1 IterSmooth: Iteration Smoothing
The mixing weight $\alpha_{t}$ increases from a small initial value (e.g., 0.1) over decoding steps toward a preset maximum(e.g., 0.2–0.4), ensuring conservative behavior early. In addition, we adopt a decode-threshold schedule that decays from 1.0 toward a preset target across steps, decoding only high-confidence positions early while progressively relaxing the criterion; which stabilizes inputs early and increases the contribution of distribution-level guidance later. Our method does not introduce new parameters or retraining, and $W_{\mathrm{emb}}$ is reused directly.
@@ B.2 Credit decoding
During decoding, we maintain a credit $C_{t}^{i,v}$ for each position $i$ and token $v\in\mathcal{V}$. This credit which quantifies how consistently a token has been favored along the generation process, serving as a temporal prior for its likelihood of being correct. Give the input $x_{t}$, let $p_{\theta}^{i}(v\mid x_{t})=\mathrm{Softmax}(f_{\theta}(x_{t})^{i}_{v})$ denote the model’s current predictive distribution, where $f_{\theta}(x_{t})$ are the logits. Let $v^{*}=\arg\max_{v}p_{\theta}^{i}(v\mid x_{t})$ be the top candidate token at position $i$. The credit is updated as follows:
@@
Here $\beta$ discounts the earlier confidece and prevents errors from accumulating and influencing future predcion. The concave transformation $(\cdot)^{\gamma}$ (with $\gamma<1$) provides relatively larger boosts to tokens with low or moderate confidence, helping correct but underconfident predictions stabilize earlier.
@@ C.1 Inference Acceleration via Trajectory Compression
While dLLMs show promise for non-sequential generation, their practical application is often hindered by high inference latency stemming from the iterative, multi-step sampling process. Inspired by the approach in Seed Diffusion Song et al. (2025), which demonstrates the value of training on high-quality generation paths, we propose a novel second-stage fine-tuning method, termed Trajectory Compression, to explicitly reduce the number of required sampling steps. The core idea is to train the model to ”jump” between non-consecutive states within an optimal generation trajectory, thereby decoding multiple tokens in a single forward pass. We refer to this resulting model as LLaDA-MoE-TD.

### v2

@@ 1 Introduction
Over the past year, diffusion-based large language models (dLLMs) have gained increasing attention in both academia and industry. Unlike conventional autoregressive (AR) models that generate tokens sequentially, dLLMs refine entire sequences in parallel through iterative denoising. This intrinsic parallelism opens new opportunities for faster decoding and enables better utilization of GPU hardware. Combined with rapid algorithmic progress, these properties make dLLMs a compelling alternative to AR LLMs. Recent work has shown that models such as LLaDA(-MoE) (Nie et al., 2025, Zhu et al., 2025) can reach performance levels comparable to strong AR baselines, including Llama (Grattafiori et al., 2024) and Qwen (Yang et al., 2024).
@@
To address these challenges, we propose dInfer, an efficient and extensible inference framework for dLLMs. dInfer modularizes inference into four components–model, diffusion iteration manager, decoding strategy, and KV-cache management–and provides well-designed APIs for flexible combinations of algorithms in each component. It supports multiple dLLM variants, including LLaDA (Nie et al., 2025), LLaDA-MoE (Zhu et al., 2025), and LLaDA-MoE-TD (Section C.1). dInfer introduces an iteration smoothing algorithm for smoother
denoising, hierarchical and credit decoding for enhanced parallel decoding, and a vicinity refresh strategy for KV-cache management to mitigate cache staleness.
@@
We evaluate inference efficiency using tokens per second (TPS) per sequence, shown in Figure 1. On HumanEval, dInfer achieves over 1,100 TPS at batch size 1, and averages more than 800 TPS across six benchmarks on a single node with $8\times$ H800 GPUs. Compared to Fast-dLLM (Wu et al., 2025), dInfer delivers more than a $10\times$ speedup while maintaining accuracy; on LLaDA-MoE it provides a $2-3\times$ speedup over QWen2.5-3B on vLLM with comparable quality.
@@ 2 Framework Design
A key advantage of dLLMs lies in their ability to perform parallel decoding within each iteration. To fully exploit this capability, advances are required in model design, diffusion iteration strategies, and decoding algorithms. At the same time, dLLMs introduce unique computational challenges. Unlike AR models–where previously computed keys and values can be cached–dLLMs employ bidirectional attention, meaning that decoding a single token can affect the representations of all tokens in the sequence. This makes straightforward KV-cache reuse infeasible and necessitates specialized cache management for efficient inference.
@@ 2.1 Diffusion Iteration Manager
dInfer currently implements two algorithms in this component. The blockwise diffusion iteration algorithm performs decoding in fixed-size spans and serves as a baseline (Algorithm 1). The iteration smoothing algorithm improves upon this by retaining token representations from the previous iteration and fusing them with the next iteration’s embeddings. This enables cross-iteration information flow, enriches contextual cues, and empirically boosts token confidence while mitigating the performance degradation typically caused by KV-cache deployment. A detailed description is provided in Appendix A.1.
@@ 2.2 Decoding Strategy
  • Threshold decoding (from Fast-dLLM (Wu et al., 2025)): commits tokens whose confidence exceeds a preset threshold.
@@ 2.3 KV-cache management
Earlier approaches introduced training-free strategies such as blockwise caching and Dual Cache (Wu et al., 2025), which reuse KV states for decoded tokens or suffixes of masked tokens. However, these methods treat cached states as static, neglecting updates from newly decoded tokens, which often degrades accuracy.
@@ 2.5 An example of the orchestration of the algorithms in dInfer
Figure 3 shows how the algorithms interact within dInfer. In each iteration, the framework tries to identify the [MASK] tokens in the active decoding block. The sequence is first embedded (potentially leveraging context from previous iterations), followed by a forward pass through the model using tensor/expert parallelism to produce logits. During the forward pass, the previous KV cache that does not hit the vicinity refresh strategy will remain unchanged and be reused. After obtaining the logits, the hierarchical/credit decoding algorithm will decide which [MASK] tokens to decode and predict their identities. In addition, the iteration smoothing algorithm will retain the logit-weighted embeddings of the sequence and incorporate them as part of the embedding for the next iteration, ensuring continuity across steps. Together, these components enable dInfer to achieve both efficiency and stability in dLLM inference.
@@ 3 Implementation Details
dInfer builds on vLLM’s backend to exploit two complementary forms of parallelism. Tensor parallelism is applied to the linear layers preceding attention modules, distributing dense computations efficiently across multiple GPUs. Expert parallelism is applied to the LLaDA-MoE model and is effective even at a batch size of 1–unlike in AR models, where expert parallelism typically requires large batch sizes. By combining tensor and expert parallelism, dInfer achieves a more than 100% improvement in inference efficiency.
@@
Diffusion iterations can suffer from CUDA stream bubbles–idle gaps of consecutive kernel launches between diffusion iterations–which waste GPU cycles and reduce throughput. To address this, dInfer applies a loop unrolling strategy that allows Python to launch CUDA kernels continuously without being blocked by stream synchronization. This reduces launch latency, keeps GPU pipelines fully occupied, and boosts iteration efficiency by about 5-10%.
We also introduce an early termination mechanism for blockwise decoding. Once an end-of-sequence (EOS) token is generated within a block, subsequent decoding steps on the remaining blocks become redundant. dInfer therefore halts the diffusion loop and fills all remaining blocks with EOS, avoiding unnecessary computation. This optimization improves inference efficiency by 15-40%.
@@ 4.1 Datasets and Configurations
All experiments are conducted on a server equipped with $8\times$ NVIDIA H800 GPUs, with PyTorch 2.9.0.dev20250831 and vLLM 0.10.1. We use a batch size of 1, a generation length of 1024 and a block size of 64 for all experiments.
@@ 4.2 Performance
As shown in Table 1, it can be clearly observed that LLaDA-MoE has comparable or higher performances than Qwen2.5-3B over six different datasets. Under the ”Without KV Cache” setting, dInfer achieves an average accuracy of 54.33, which is higher than Fast-dLLM and is comparable to the performance of QWen2.5-3B in vLLM and LLaDA-MoE reported in its paper (Zhu et al., 2025), while achieving $6.5\times$ speedup over Fast-dLLM. When KV cache is enabled, dInfer achieves higher accuracy than Fast-dLLM (53.96 vs. 52.15) and delivers $6\times$ speedup over Fast-dLLM. When achieving similar model performance, dInfer achieves over $10\times$ speedup (a TPS of 680.71) over Fast-dLLM (a TPS of 63.61) across the six benchmarks. dInfer is also $2.5\times$ faster than Qwen-2.5 3B (a TPS of 277.45) in vLLMs.
@@ 5 Conclusion
In this work, we present dInfer, an efficient and extensible inference framework for dLLMs. By decomposing inference into modular components and incorporating optimizations such as hierarchical decoding, credit decoding, iteration smoothing, and vicinity KV-cache refresh, dInfer effectively addresses the key bottlenecks of high computation cost and limited parallel decoding efficiency of dLLMs. Extensive experiments on LLaDA-MoE demonstrate that dInfer achieves state-of-the-art throughput–exceeding 1,100 TPS on $8\timesH$800 GPUs–while maintaining output quality. We believe that dInfer provides both a practical toolkit and a standardized platform to accelerate research and development in the rapidly growing field of dLLMs.
@@ A.1 IterSmooth: Iteration Smoothing
The mixing weight $\alpha_{t}$ increases from a small initial value (e.g., 0.1) over decoding steps toward a preset maximum (e.g., 0.2–0.4), ensuring conservative behavior early. In addition, we adopt a decode-threshold schedule that decays from 1.0 toward a preset target across steps, decoding only high-confidence positions early while progressively relaxing the criterion, which stabilizes inputs early and increases the contribution of distribution-level guidance later. Our method does not introduce new parameters or retraining, and $W_{\mathrm{emb}}$ is reused directly.
@@ B.2 Credit decoding
During decoding, we maintain a credit $C_{t}^{i,v}$ for each position $i$ and token $v\in\mathcal{V}$. This credit which quantifies how consistently a token has been favored along the generation process, serving as a temporal prior for its likelihood of being correct. Given the input $x_{t}$, let $p_{\theta}^{i}(v\mid x_{t})=\mathrm{Softmax}(f_{\theta}(x_{t})^{i}_{v})$ denote the model’s current predictive distribution, where $f_{\theta}(x_{t})$ are the logits. Let $v^{*}=\arg\max_{v}p_{\theta}^{i}(v\mid x_{t})$ be the top candidate token at position $i$. The credit is updated as follows:
@@
Here $\beta$ discounts the earlier confidence and prevents errors from accumulating and influencing future prediction. The concave transformation $(\cdot)^{\gamma}$ (with $\gamma<1$) provides relatively larger boosts to tokens with low or moderate confidence, helping correct but underconfident predictions stabilize earlier.
@@ C.1 Inference Acceleration via Trajectory Compression
While dLLMs show promise for non-sequential generation, their practical application is often hindered by high inference latency stemming from the iterative, multi-step sampling process. Inspired by the approach in Seed Diffusion (Song et al., 2025), which demonstrates the value of training on high-quality generation paths, we propose a novel second-stage fine-tuning method, termed Trajectory Compression, to explicitly reduce the number of required sampling steps. The core idea is to train the model to ”jump” between non-consecutive states within an optimal generation trajectory, thereby decoding multiple tokens in a single forward pass. We refer to this resulting model as LLaDA-MoE-TD.